OpenStudy (anonymous):

NEED HELP ASAP!!! What is the missing coefficient in the equations given that p(3)=567? p(x)=_______x^4

OpenStudy (perl):

p(x) = A * x^4 567 = A*(3)^4 solve for A A = 567 / 3^4

OpenStudy (anonymous):

So the answer is 567/3^4?

OpenStudy (perl):

yes and that equals 7

OpenStudy (anonymous):

Thank you!